### v1

# A Novel Passive Occupational Shoulder Exoskeleton With Adjustable Peak Assistive Torque Angle For Overhead Tasks

Abstract
[…]
{IEEEkeywords}
Shoulder exoskeletons, musculoskeletal disorders, overhead work, efficient torque generator, ergonomics.
00footnotetext: Copyright (c) 2021 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
1 Introduction
[…]
OSEs can be classified into active, semi-active, and passive categories. In general, active OSEs can provide higher, more precise, and versatile assistive torques [8, 9, 10]. However, they have larger volumes and higher maintenance costs [11]. As an alternative solution, semi-active OSEs can automatically adjust torque profiles by employing small low-power motors [12], such as the H-PULSE [9], [13]. Compared to active ones, it is more compact and has lower energy consumption while maintaining effectiveness. However, compared to passive ones, the maintenance and manufacturing costs of semi-active OSEs are higher, which would limit their widespread adoption and application. Passive OSEs (POSEs) are characterized by their low cost and lightweight design [14], [15], which provide arm gravity compensation during overhead works without external power sources. Recent studies have demonstrated that wearing POSEs can reduce muscle activation during overhead tasks performed in laboratory and field settings [16, 17, 18]. Even though POSEs have been commercialized, some pending issues remain, thus impeding user acceptance and assistive efficiency [19].
One of the main issues is the insufficient kinematic compatibility between the exoskeleton and the user [20]. POSEs can be classified based on the presence or absence of an exoskeletal shoulder structure above the human shoulder [11]. Exoskeletons with a rigid structure above the human shoulder, such as the ShoulderX [21] and Airframe [22], typically feature a degree of freedom (DoF) next to the shoulder to follow shoulder flexion/extension (F/E) movements. These exoskeletons are compact and thus are suitable for performing overhead tasks in narrow spaces. However, they may induce collisions between the shoulder and the exoskeleton at high elevation angles, resulting in an insufficient range of motion (ROM). Exoskeletons without a structure above the shoulder, such as the SPM [23] and Exo4Work [24], avoid collisions at high elevation angles.However, they generally incorporate multi-DoF structures in the back to accommodate shoulder movements, which leads to a larger volume. The drawback may result in collisions with the surrounding objects in narrow spaces. In existing exoskeletons, achieving both sufficient ROM and compactness is mutually exclusive. To achieve a balance between them, Junsoo Kim et al. [11] designed an exoskeleton with tilted shoulder structures for assisting overhead tasks. However, it still cannot provide a sufficient ROM for horizontal F/E movements, which may induce a decreased user experience. There is an urgent need for an exoskeleton that can simultaneously provide sufficient ROM and compactness.
Figure 1: The torque profiles when there is a match (orange line) or mismatch (blue line) between the PATA and the target angle of the task in an overhead task with a target angle of 120º. Angle difference refers to the discrepancy between the PATA and target angle of the task.
Another main issue is the task-specific or individual-specific design, corresponding to a fixed assistive torque profile and thus a fixed assistive region (i.e. the subset of ROM where the exoskeleton can assist), which may not be optimal for other tasks [25], [26]. Overhead tasks relate to different positions, especially different elevation angles of the shoulder (e.g., screwing task of the automobile manufacturing industry [21, 27, 28], aircraft squeeze riveting and sealing tasks [29], [30], the welding task of shipbuilding [11], order picking task [16] and so on). Consequently, various overhead tasks require different optimal torque profiles [13], [24]. In addition, the same task may have different target angles for individuals of different heights, which requires providing personalized torque profiles [31]. The torque profile should be adaptive to tasks and subjects, essentially characterized by the amplitude and peak assistive torque angle (PATA). The amplitude of the peak assistive torque can be easily tuned by adjusting the extension length of the spring. For example, the MATE can manually change the peak torque by a knob [32], and H-PULSE can adjust the spring length by a spindle drive coupled with a servomotor [9], [13]. The timing of the peak assistive torque ensures that the POSE can provide timely assistance when the torque profile reaches its peak value at specific elevation angles of the shoulder [19]. Fig. 1 depicts the torque profiles when there is a match or mismatch between the PATA and the target angle of the task in an overhead task with a target angle of 120º. When there is a match between them, it ensures timely assistance and reduces energy loss, making it a more efficient assistive strategy. However, current POSEs can only change the PATA through reconstruction, which means they cannot provide torque profiles for different tasks or individuals, i.e., insufficient adaptability [26], [31]. In summary, different tasks and individuals should have different PATAs, lacking in current POSEs. Therefore, an adjustable PATA is needed to enhance task and subject adaptability.
The object of this study is to develop, implement, and preliminarily evaluate a novel POSE for overhead tasks, named HIT-POSE. Our prototype can provide a sufficient ROM of the human shoulder and possess an adjustable PATA, thus adapting to various overhead tasks and individuals.The major contributions of our work are as follows:
(a) We designed a torque generator with adjustable PATA to adapt to different overhead tasks and individuals. To the best of our knowledge, it’s the first torque generator that can provide adjustable PATA to achieve more efficient assistance.
(b) We developed an ergonomic shoulder structure through kinematic modeling and simulation, which can guarantee sufficient ROM for the shoulder while remaining compact (without redundant structures in the back), thus reducing the risk of collisions with the environment.
(c) We also investigated how the modulation of the PATA could affect objective and subjective evaluation indicators in overhead tasks via experimental assessment, thus validating the effectiveness of the exoskeleton.
2 System Design of HIT-POSE
2.1 System Overview
Overhead work is defined as working with hands above acromion or over 60° shoulder flexion or abduction [33]. Therefore, we can define the workspace for overhead tasks as the sagittal flexion of the shoulder with 60° to 168.8° (the maximum shoulder flexion angle for healthy adult male [34]), abduction angle of the shoulder greater than 60°, and horizontal flexion of the shoulder with 0° to 162.2° (the maximum shoulder horizontal flexion angle for healthy adult male [11]). In addition, the total weight of the exoskeleton should be less than 5kg [19].
HIT-POSE is designed to compensate for shoulder gravitational torque while providing sufficient ROM for the shoulder, as illustrated in Fig. 2. The total weight of the device 2.8 kg. It is suitable for people with a back length of 43.6-66.6 cm and a shoulder width of 29.6-45.6 cm, and the volume of our device is 22cm × 48cm × 15cm. HIT-POSE comprises four main modules: (a) the ergonomic shoulder structure, (b) the torque generator with adjustable PATA, (c) the physical user-exoskeleton interface (PUEI), and (d) the size regulation module.
Figure 2: An overview of HIT-POSE.
2.2 Ergonomic Shoulder Structure
We conducted kinematic modeling of the exoskeleton and the user and simulated ROM to select optimal design parameters. The proposed exoskeleton has a horizontal F/E joint above the shoulder and lacks redundant structures in the back. Our object here is to provide sufficient ROM for the shoulder, especially at high elevation angles.
The intersection of the rotational axes of the sagittal and horizontal F/E joints is considered the center of the shoulder of the exoskeleton (CSE), as shown in Fig. 3. The CSE is close to the center of the shoulder of the user (CSU), enabling the torque generator to trace the shoulder movements with a relatively low DoF (Fig. 3). However, it does not coincide with the CSU, indicating the necessity to find the optimal design parameters. As explained in [35] and [36], the shoulder motion can be observed with humeral elevation and posterior displacement of the CSU.
Fig. 3 demonstrates the extracted design parameters ($\phi$, $d_{v}$, $d_{b}$), which can reconfigure the shoulder structure. We defined the $\phi$-plane as the plane containing the front-and-rear adjustable linkage and perpendicular to the horizontal plane (Fig. 3). Viewed from the $\phi$-plane, the pitch angle ($\phi$) is the angle between the front-and-rear adjustable linkage and horizontal plane, and the vertical distance ($d_{v}$) refers to the perpendicular distance from the CSU to the front-and-rear adjustable linkage, as presented in Fig. 3a. Bias distance $d_{b}$ is defined as the distance between the CSE and CSU from the top view (Fig. 3b). $\phi$ and $d_{v}$ can determine the configuration of the front-and-rear adjustable linkage and shoulder bending linkage, respectively. $d_{b}$ can be regulated by adjusting the length of the front-and-rear adjustable linkage, to adapt to different users and ensure comfort. The orange clear orb (Fig. 3a) can be enlarged by increasing $\phi$ and $d_{v}$, provided for humeral elevation and avoiding collision at high elevation angles. Increasing $d_{b}$ shifts CSE posteriorly, compensating for the displacement of the CSU during shoulder movement. Overall, a proper combination of design parameters ($\phi$, $d_{v}$, $d_{b}$) can increase the user’s ROM when wearing the exoskeleton.
Figure 3: Definition of the design parameters ($\phi$, $d_{v}$, $d_{b}$). (a) $\phi$ and $d_{v}$ are defined from the $\phi$-plane view, the orange clear orb represents the room between the CSU and the front-and-rear adjustable linkage. (b) $d_{b}$ is defined from the top view.
Figure 4: Kinematic models. (a) The user. (b) The exoskeleton.
Table 1: DH PARAMETERS OF THE USER
$\alpha_{i-1}$ $a_{i-1}$ $d_{i}$ $\theta_{i}$
1’($\theta_{h,user}$) 0 0 0 $\pi$/2+$\theta_{h,user}$
2’($\theta_{s,user}$) -$\pi$/2 0 0 -$\pi$/2+$\theta_{s,user}$
Arm sleeve 0 $D_{user}$ 0 0
Table 2: DH PARAMETERS OF THE EXOSKELETON
$\alpha_{i-1}$ $a_{i-1}$ $d_{i}$ $\theta_{i}$
1($d_{b}$) $\pi$/2 0 $d_{b}$ 0
2($\phi$, $d_{v}$) -$\pi$/2+$\phi$ 0 $d_{v}$ 0
3($\theta_{h,exo}$) 0 0 0 $\pi$/2+$\theta_{h,user}$
4($d_{v}$) 0 0 -$d_{v}$ 0
5($\theta_{s,exo}$) -$\pi$/2 0 0 -$\phi$-$\pi$/2+$\theta_{s,exo}$
Arm sleeve 0 $D_{exo}$ 0 0
Simulation $\mathrm{I}$ Calculate the user’s ROM when wearing the exoskeleton
Input $\theta_{h,exo}$, $\theta_{s,exo}$, $\phi$, $d_{v}$, $d_{b}$, $D_{exo}$, $D_{user}$
Output ROM simulation results
1: Calculating the position of the arm sleeve ($p_{x}$, $p_{y}$, $p_{z}$) via forward kinematics of the exoskeleton;
2: Solving $\theta_{h,user}$ and $\theta_{s,user}$ using inverse kinematics of the user and position of arm sleeve ($p_{x}$, $p_{y}$, $p_{z}$);
3: Checking collision between the exoskeleton and user;
4: Obtain the non-collisional joint angles of the user.
To better explore the effect of design parameters on ROM, the kinematic analyses were conducted, as shown in Fig. 4. In this study, we only consider 2-DoFs of the user’s shoulder, a horizontal and sagittal F/E DoF, whose angles were denoted by $\theta_{h,user}$ and $\theta_{s,user}$ (Fig. 4a). Similarly, the exoskeleton’s shoulder has 2-DoFs, whose angles are denoted by $\theta_{h,exo}$ and $\theta_{s,exo}$ (Fig. 4b). These four angles represent the ROM for both the human and the exoskeleton. $D_{user}$ and $D_{exo}$ represent the distances from the CSU and CSE to the arm sleeve (Fig. 4), respectively.
Tables 1 and 2 demonstrate the specific Denavit-Hartenberg (DH) parameters of both models. The association between the two models is that the coordinate origin (the CSU in Fig. 4) and arm sleeve are in the same location. Based on the characteristics, we formulated a simulation pipeline via the analytical method to calculate the user’s ROM when wearing the exoskeleton, as displayed in Simulation $\mathrm{I}$. Notably, the principle for checking collisions is that the angles of the user cannot be greater than that of the exoskeleton.
A previous study has stated that the CSU is about 2 cm below the acromion [37]. Additionally, the CSU is vertically displaced by around 6 cm during shoulder movement [38]. Such prior knowledge provides a basis for choosing the value of $d_{v}$. We provide seven sets of design parameters and their ROM simulation results, as shown in Fig. 5. The area below each set of curves is the corresponding ROM. When $\theta_{h,user}$=0, the maximum value of $\theta_{s,user}$ is obtained. The ROM can be enlarged by increasing $\phi$ and $d_{v}$, thus preventing collisions at high elevation angles. To maximize the user’s ROM, $\phi$ and $d_{v}$ were set to 15º and 80 mm, respectively. As $d_{b}$ increases, the CSE moves toward the posterior of the user, causing increased ROM, but too large $d_{b}$ will lead to discomfort. Eventually, we take the penultimate set ($\phi$=15°, $d_{v}$=80mm, $d_{b}$=10mm) as the final design parameters. In this set, the ROM is sufficient to ensure the user’s natural movement and complete overhead tasks.
Figure 5: ROM simulation results of the user with seven sets of design parameters ($\phi$, $d_{v}$, $d_{b}$).
Figure 6: The overall structure and design parameters of the torque generator with adjustable PATA. (a) Components of the torque generator. (b) Torque generator with parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$) after the wire rope is attached to the sagittal F/E joint bar (Phase $\mathrm{I}$). (c) Torque generator with parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$) before the wire rope is attached to the sagittal F/E joint bar (Phase $\mathrm{II}$). 1: Sagittal F/E joint bar. 2: Pulley fixed to the shoulder bending linkage. 3-6: Pulleys fixed to the carbon fiber shell. 7: Parallel spring group. 8: 8-shaped loop. 9: Spring pretension tuning module. $L_{exo}$: the force arm of the assistive torque.
Figure 7: The principal of the assistive torque profile. (a) Overlapped location configurations of the force arm $L_{exo}$, $\theta_{c}$ is the critical angle between phases $\mathrm{I}$ and $\mathrm{II}$, $r_{2}$ is the distance between the guide rail and the sagittal F/E joint, $\beta$ is the angle of pulley3, as shown in Fig. 6b Torque profile $\tau_{exo}$ along with the force arm $L_{exo}$ and total extension length of the parallel spring group $\Delta{L}$. The area shaded in light orange is Phase $\mathrm{I}$, and that in light blue is Phase $\mathrm{II}$.
Figure 8: The relationship between $\beta$ and PATA.
2.3 Torque Generator with Adjustable PATA
The torque generator is composed of a sagittal F/E joint bar, PATA adjustable module, parallel spring group, spring pretension tuning module, carbon fiber shell, and five pulleys with inner bearings (Fig. 6a). A wire rope of 1mm diameter connects the parallel spring group to pulley 2. The pulleys 3 and 4 can be moved vertically through the guide rail, resulting in the adjustable PATA. The 8-shaped loop with a rotational DoF connects the parallel spring group with the spring pretension tuning module. It ensures that the parallel spring group does not experience axial rotation when screwing the spring pretension tuning module. Three identical springs (i.e., $L_{s}$=$L_{AB}$=$L_{CD}$=$L_{EF}$) are installed in the parallel spring group to obtain a high energy storage capability.
The assistive principle is to store or release the elastic potential energy of springs as the arm is lowered or raised, respectively. Since the torque generator follows the arm movement via an arm sleeve, it can be considered that the sagittal F/E angle of the exoskeleton ($\theta_{s,exo}$) is approximately equal to that of the user ($\theta_{s,user}$). The generated assistive torque ($\tau_{exo}$) can be formulated as:
$\tau_{exo}=F_{s}L_{exo}.$ (1)
where $F_{s}$ is the tension generated by springs, and $L_{exo}$ denotes the perpendicular distance between point O and the wire rope (Fig. 7a). Furthermore, $F_{s}$ can be expressed as:
$F_{s}=K\Delta{L}.$ (2)
[…]
$\Delta{L}(\theta_{s,exo})=L_{A^{\prime}B^{\prime}}(\theta_{s,exo})-L_{A^{\prime}B^{\prime}}(\max(\theta_{s,exo}))+L_{i}.$ (3)
where $L_{A^{\prime}B^{\prime}}$ is the length of the wire rope routed from A’ to B’ (Fig. 6b and 6c). A’ and B’ are the starting and ending points of the wire rope that attach to pulley 2 and pulley 4, respectively, and $L_{i}$ denotes the initial extension length caused by the spring pretension tuning module. The minimum value of $L_{A^{\prime}B^{\prime}}$ is obtained at the maximum value of $\theta_{s,exo}$, which is set to 170º due to the safety limit.
Based on (1) to (3) and Fig. 7b, the main factor affecting PATA is $L_{exo}$, rather than $\Delta{L}$, as it is monotonically decreasing. The primary parameters affecting the phase shift of $L_{exo}$ are $\alpha$ and $\beta$, denoting the angles of pulleys 2 and 3, as shown in Fig. 6b and 6c. When $\alpha$ decreases or $\beta$ increases, PATA would decrease and thus the phase of $L_{exo}$ curve would shift to the left (Fig. 7a and 7b). The assistive torque would be influenced by $L_{exo}$ and also have a phase shift (as shown in Fig.1). To better investigate the trend of PATA, when $\alpha$=$\beta$=0, we defined the PATA as a reference, denoted by $\theta_{0}$. The PATA ($\theta_{PATA}$) can be derived as:
$\theta_{PATA}=\theta_{0}+\alpha-k\beta.$ (4)
$\theta_{0}=90\textdegree-arctan(\frac{r}{r_{2}}).$ (5)
where r is the radius of pulley 3, which is 7mm in this paper, thus $\theta_{0}$ is 80º; and k is a variable positive coefficient around 1, due to the nonlinear effect of $\Delta{L}$ on $\tau_{exo}$. To change the PATA for adapting different overhead tasks, we adjusted $\beta$ by moving pulleys 3 and 4 along the guide rail, the relationship between $\beta$ and PATA can be found in Fig. 8. We employed $\alpha$ to control the maximum sagittal F/E joint angle ($\theta_{s,exo}$). $\alpha$ can be adjusted by changing the mounting position of pulley 2 through the pre-drilled holes, and its mappings to the maximum of $\theta_{s,exo}$ are indicated by the dotted line in Fig. 9e.
2.4 Parameters Impact on Torque Profiles
In this subsection, we analyzed the impact of design parameters on the peak value and PATA of the torque profile and conducted a simulation of torque profiles. Table 3 shows the parameters of the torque generator.
As visualized in Fig. 6b and 6c, the generated assistive torque can be divided into two phases based on whether the wire rope is routed to the sagittal F/E joint bar. When the wire rope is attached to the sagittal F/E joint bar (Phase $\mathrm{I}$, Fig. 6b), the assistive torque remains low to avoid hindering normal activities like walking, corresponding to the low assistance zone (Fig. 7b). In contrast, when the wire rope is detached from the sagittal F/E joint bar (Phase $\mathrm{II}$, Fig. 6c), the assistive torque increases rapidly to provide effective assistance for the target task, corresponding to the high assistance zone (Fig. 7b). We defined the critical angle for switching between phases $\mathrm{I}$ and $\mathrm{II}$ as $\theta_{c}$ (Fig. 7a), which is dependent on $\alpha$ and $\beta$ (Fig. 6b and 6c). As $\alpha$ increases or $\beta$ decreases, $\theta_{c}$ increases (Fig. 9e - 9f). The wire rope coupled with the sagittal F/E joint bar formulates a passive clutch mechanism that allows for switching between low and high assistance zones. If there is no sagittal F/E joint bar (i.e., orange dotted line in Fig. 6b), the torque generator will pull the arm backward, which poses a safety hazard to the user. It is worth noting that this mechanism also provides a torque of approximately 2 N·m, although it is much smaller than the peak assist torque, it may still reduce the user’s acceptance of the exoskeleton.
Table 3: THE EFFECTS OF PARAMETERS ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) ON TORQUE ASSISTIVE AND PATA IN BOTH PHASES
Requirement Parameter Value
Peak torque K 12N/mm
$L_{i}$ 30mm
$r_{1}$ 20mm
$r_{2}$ 40mm
Torque at phase $\mathrm{I}$ $r_{3}$ 3mm
Critical angle $\alpha$ 40°
Table 4: THE EFFECTS OF PARAMETERS ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) ON TORQUE ASSISTIVE AND PATA IN BOTH PHASES
Parameter Phase $\mathrm{I}$ Phase $\mathrm{II}$
$\tau_{exo}$ $\tau_{exo}$ PATA
$\alpha$ $\mathrm{NI^{a}}$ $\mathrm{N^{b}}$ $\mathrm{P^{c}}$
$\beta$ NI P N
$r_{1}$ NI P NI
$r_{2}$ NI NI NI
$r_{3}$ P NI NI
$L_{i}$ P P NI
a: No Impact; b: Negative; c: Positive.
Figure 9: The simulated assistive torque ($\tau_{exo}$) profiles were obtained via tuning of the assistive parameters. (a) $L_{i}$, (b) $r_{1}$, (c) $r_{2}$, (d) $r_{3}$, (e) $\alpha$, (f) $\beta$.
$L_{exo}$ and $L_{A^{\prime}B^{\prime}}$ can be geometrically calculated in the two phases. Based on (1) to (3) , the peak value of the assistive torque increases with the stiffness (K) and initial extension length ($L_{i}$) of the parallel spring group in both phases. In phase $\mathrm{I}$, the primary parameter affecting $L_{exo}$ and $L_{A^{\prime}B^{\prime}}$ is $r_{3}$, the radius of the sagittal F/E joint bar. As $r_{3}$ increases, $L_{exo}$ and $L_{A^{\prime}B^{\prime}}$ increase, resulting in an increase in assistive torque within the low assistance zone. In phase $\mathrm{II}$, the primary parameters affecting $L_{exo}$ and $L_{A^{\prime}B^{\prime}}$ are $\alpha$, $\beta$, $r_{1}$, where $r_{1}$ represents the distance between pulley 2 and the sagittal F/E joint. As $r_{1}$ increases, $L_{exo}$ and $L_{A^{\prime}B^{\prime}}$ increase, causing an increase in the peak value of the assistive torque. When $\alpha$ decreases or $\beta$ increases, $L_{A^{\prime}B^{\prime}}$ increases, leading to an increase in the peak value of the assistive torque, and the impact on $L_{exo}$ has been stated in subsection C. We also extracted and tested $r_{2}$, the distance between the guide rail and the sagittal F/E joint, as a design parameter. The result showed that it essentially has no significant impact on the peak or phase of the assistive torque, allowing us to choose an appropriate value according to the requirements of the design space. Table 4 summarizes the effects of the parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) on the assistive torque and PATA.
Fig. 9 demonstrates how the parameters affect the assistive torque profiles through simulations with an ideal situation (i.e., no deformation and friction). The simulation results in Fig. 9 are in accord with our previous analyses. In particular, the parameters $L_{i}$, $\alpha$ and $\beta$ can be easily adjusted, and K can be easily changed by replacing the spring. Consequently, this implies that the main characteristics of the torque profile can be tuned by reconfiguring the exoskeleton (i.e. by tuning the pre-tension of springs or the mounting position of pulley 2), rather than reconstructing it. As far as we know, this is the first exoskeleton that can simultaneously modify these characteristics of the torque profile without the need of reconstruction.
2.5 PUEI and Size Regulation Module
The PUEI contains a lower back linkage and all components that attach the exoskeleton to the user, e.g., shoulder straps, chest buckles, and waist band. These flexible fabrics ensure a compliant human-exoskeleton interface. Furthermore, the lower back linkage is made of aluminum alloy, which transmits the force compensated by the torque generator to the waist.
The size regulation module is designed to accommodate diverse ergonomic needs while ensuring both comfort and safety. The height and width of the back are adjustable through upper and lower back linkages, as well as lateral back slides. Additionally, the flexible fabric allows for precise tightening of the exoskeleton around the user’s waist, back, and arms.
3 EXPERIMENTAL EVALUATIONS
Two experimental sessions were conducted to verify the HIT-POSE has (a) a sufficient ROM (session 1), and (b) an ability and adaptability to assist different overhead tasks (session 2).
3.1 Participants
Two separate groups of participants were recruited for this study. Five healthy subjects participated in session 1 (all male, all right-handed, age: 24.1 ± 1.64 years, height: 173.8 ± 3.27 cm, weight: 70 ± 9.06 kg), and ten healthy subjects participated in session 2 (all male, all right-handed, age: 25.1 ± 2.0 years, height: 177 ± 4.3 cm, weight: 68.3 ± 9.3 kg). All participants had no prior experience with the exoskeleton and signed an informed consent before the experiment. The experimental protocol received approval from the Chinese Ethics Committee of Registering Clinical Trials (ChiECRCT20200319).
3.2 Data Recordings
ROM of the right-side shoulder was measured by the motion capture system (100Hz, VICON, Oxford Metrics, Oxford, UK). Muscular activities were monitored using the surface EMG electrodes ($1$Hz, Delsys Inc., Natick, MA, USA). Meanwhile, acceleration signals were recorded through an IMU (148Hz, Delsys Inc., Natick, MA, USA). Notably, the motion capture system did not work simultaneously with EMG. Thus, we did not synchronize them through a trigger.
Figure 10: Experimental tasks. (a) Shoulder movements in session 1. (b) Screwing task in session 2.
Figure 11: The definition of the low (left) and high (right) heights.
3.3 Experimental Tasks
In session 1, participants were directed to maximize sagittal and horizontal F/E movements, at a self-selected constant pace, as shown in Fig. 10a. In session 2, participants were guided to screw the M6 bolts with a hex key (mass: 8.6 g), as shown in Fig. 10b. Nuts were embedded in an aluminum frame, which was set to two heights: low height was defined as hand with the shoulder at about 90º, and high height was defined as hand with the shoulder at about 120º (Fig. 11).
3.4 Procedure
Before the experiment, participants were introduced to the experimental paradigm and familiarized with HIT-POSE and tasks. In session 1, tasks were performed under two conditions: without the exoskeleton (NO HIT-POSE) and with the exoskeleton (HIT-POSE). Each participant completed six trials: three in each condition, with each trial consisting of five movements and lasting about 30 seconds. In session 2, the exoskeleton was configured for screwing at low and high heights (Fig. 12). Both configurations had the same peak torque but different PATAs for the respective heights. A match between the PATA and the task angle was defined as HIT-POSE Match; otherwise, it was HIT-POSE Mismatch. Each task was performed under three conditions: NO HIT-POSE, HIT-POSE Match, and HIT-POSE Mismatch (Fig. 12), with each trial lasting 2 minutes per condition.
Surface EMG electrodes were placed unilaterally on the right-side muscles according to SENIAM guidelines [39]: anterior deltoid (AD), middle deltoid (MD), posterior deltoid (PD), bicep brachii (BB), pectoralis major (PM), trapezius (TR), latissimus dorsi (LD), erector spinae (ES). Additionally, an IMU was placed on the right wrist to measure the acceleration. Then, the EMG signals were normalized for each participant and task using the maximal amplitude of each muscle during NO HIT-POSE conditions [40, 41, 42]. The order of conditions and tasks was randomized across participants to avoid order effects. After each trial, participants completed the NASA-TLX [43] to assess perceived workload (PW) and the Borg CR-10 [44] for rating perceived exertion (RPE). After all exoskeleton trials, they completed the System Usability Scale (SUS) [45]. Fig. 12 presents the experimental procedure for session 2.
Figure 12: The experimental flowchart (left) in session 2 and procedure in one condition (right). The order of the conditions and tasks is randomized.
3.5 Data and Statistical Analysis
In session 1, mean values of the maximum sagittal and horizontal F/E angles were calculated using shoulder angles measured by VICON and corresponding hand workspace were calculated using the position of the reflective point on hand measured by VICON among all participants in NO HIT-POSE and HIT-POSE conditions. In session 2, acceleration signals were resampled to $1$Hz, and variances were calculated using a 30ms sliding window. A set variance threshold determined the start and end times of each trial. EMG signals were processed to compute the linear envelopes by setting them to zero-mean, applying band-pass filter (4th, 20-350Hz), full-wave rectification, low-pass filter (6Hz), and normalization. For each trial, the root mean square (RMS) of each muscle was obtained from the linear envelope via a 30ms time sliding window during the phase of task execution.
Statistical analyses were used to assess the effectiveness of the exoskeleton in reducing muscle activations and subjective evaluation indicators (i.e., RPE and PW scores). The Shapiro-Wilk test was employed to check data normality. If the normal distribution was not satisfied, the non-parametric analysis of variance (ANOVA) was conducted to check for across-condition differences, otherwise, the parametric ANOVA was applied. A post-hoc analysis was conducted when a significant effect was observed. Statistical significance was concluded when p $<$ 0.05.
4 RESULTS
4.1 ROM
Table 5 reports measured ROM averaged among all participants. No significant differences were observed between conditions. The absolute error between the maximum value (169.71°) of the penultimate set in the simulation of Fig. 5 and the maximum sagittal F/E angle (164.46°) in HIT-POSE condition of the experiment was 5.25°. Fig. 13 shows the averaged hand workspace (normalized by the arm length of each subject) of the maximum sagittal and horizontal F/E angles among all participants in both conditions. In both movements, the coordinate origin is the hand’s position in the neutral posture before movements. The positive x-axis points to the right side of the body, the positive y-axis points forward, and the positive z-axis points upward. There was no significant statistical difference between conditions for both tasks.
Figure 13: Averaged hand workspace (normalized by the arm length of each subject) of the maximum sagittal and horizontal F/E angles among all participants in both conditions. (a) Hand workspace of sagittal F/E, (b) Hand workspace of horizontal F/E. The error bar represents the mean ± standard deviation.
Table 5: MEASURED ROM WITHOUT AND WITH HIT-POSE
Condition Maximum sagittal F/E angle (deg) Maximum horizontal F/E angle (deg)
NO HIT-POSE 165.13 ± 2.15 159.74 ± 6.27
HIT-POSE 164.46 ± 2.21 158.28 ± 7.20
4.2 Muscular Activations
For all monitored muscles, statistically significant interactions were found in both conditions, as shown in Table 6. Fig. 14 demonstrates the results of muscular activations from pairwise comparisons. In HIT-POSE Match conditions for both tasks, significantly decreased muscular activations in all monitored muscles were found compared to NO HIT-POSE condition. For screwing at a low height, the muscular activations of MD, PD, BB, PM, and LD muscles in HIT-POSE Match condition were significantly reduced compared to that in HIT-POSE Mismatch condition; in addition, the muscular activation of the BB muscle in HIT-POSE Mismatch condition was also significantly reduced compared to that in NO HIT-POSE condition. For screwing at a high height, AD, MD, PD, and PM muscles in HIT-POSE Match condition exhibited significantly decreased EMG activations compared to that in HIT-POSE Mismatch condition; moreover, BB and TR muscles in HIT-POSE Mismatch condition exhibited significantly decreased EMG activations compared to that in NO HIT-POSE condition.
[…]
Table 6: F AND P-VALUES OF THE ANOVAS ON THE MUSCULAR ACTIVATIONS DURING THE TWO TASKS.
Muscles Screwing at a low height Screwing at a high height
F-value P-value F-value P-value
AD 11.65 0.0006 13.03 0.0003
MD 16.84 0.0001 12.33 0.0004
PD 18.75 0.0000 19.91 0.0000
BB 13.75 0.0002 13.25 0.0003
PM 11.23 0.0007 25.60 0.0000
TR 10.08 0.0012 11.85 0.0005
LD 10.69 0.0009 8.10 0.0031
ES 12.00 0.0005 11.61 0.0006
[…]
Muscles Screwing at a low height Screwing at a high height
[…]
AD 22.9||||48.3 11.5||||24.3 23.8||||48.5 10.6||||21.6
MD 15.6||||30.3 6.5||||12.5 19.5||||37.9 8.2||||15.8
PD 18.5||||35.6 7.0||||13.5 23.3||||44.7 8.9||||17.1
BB 15.5||||41.9 7.6||||20.6 20.1||||49.6 11.3||||27.7
PM 19.0||||43.4 5.5||||12.5 17.9||||31.9 5.4||||9.6
TR 19.6||||40.9 8.8||||18.2 20.2||||42.2 11.3||||23.6
LD 17.4||||47.0 7.6||||20.6 15.7||||42.5 7.2||||19.6
ES 15.9||||43.3 8.0||||21.8 17.0||||43.6 8.3||||21.4
A||||B: A and B denote absolute and relative percentage reduction with respect to NO HIT-POSE condition, respectively.
Figure 14: Muscular activations within different conditions for two tasks from pairwise comparisons. (a) Screwing at a low height. (b) Screwing at a high height. The error bar represents the mean ± standard deviation. *: p $<$ 0.05; **: p $<$ 0.01; ***: p $<$ 0.001.
4.3 Subjective Feedback
For RPE scores, statistically significant interactions were found in both conditions (F = 14.58, p = 0.0002 for low height; F = 10.55, p = 0.0009 for high height). Fig. 15 shows the RPE scores averaged across participants from pairwise comparisons. In HIT-POSE Match conditions for both tasks, RPE scores showed a significant difference relative to the other two conditions. For physical demand, effort, and frustration aspects of NASA-TLX, statistically significant interactions were found in both conditions (Physical Demand: F = 5.52, p = 0.0135 for low height; F = 5.08, p = 0.0179 for high height. Effort: F = 7.76, p = 0.0037 for low height; F = 6.85, p = 0.0061 for high height. Frustration: F = 6.34, p = 0.0082 for low height; F = 4.08, p = 0.0345 for high height). For both tasks in HIT-POSE Match conditions, significant differences were found in physical demand, effort, and frustration compared to NO HIT-POSE conditions (Fig. 16). In addition, the average SUS score was 79.7 ± 5.9.
Figure 15: RPE scores for each task among the three conditions from pairwise comparisons. The error bar represents the mean ± standard deviation. *: p $<$ 0.05; **: p $<$ 0.01; ***: p $<$ 0.001.
Figure 16: PW scores for each task among the three conditions from pairwise comparisons. (a) Screwing at a low height. (b) Screwing at a high height. MD = mental demand, PD = physical demand, TD = temporal demand, P = performance, E = effort, F = frustration. The error bar represents the mean ± standard deviation. *: p $<$ 0.05; **: p $<$ 0.01; ***: p $<$ 0.001.
5 DISCUSSION
5.1 Kinematics Compatibility
HIT-POSE is equipped with an ergonomic shoulder structure to ensure sufficient ROM for the user. This implies favorable kinematics compatibility, which has a profound impact on the exoskeleton’s usability and user acceptance. Typically, some passive shoulder exoskeletons integrate a multi-DoF chain or simplified two-link hinge to increase the ROM, such as the SPM [23] and HUST-EV [46]. However, they may collide with nearby objects easily due to the large volume and protrusions. In addition, the design parameters were fixed, thus lacking reconfigurability. This reduces the adaptability and increases the manufacturing costs of the exoskeleton.
To ensure sufficient ROM while compactness, we extracted three parameters ($\phi$, $d_{v}$, $d_{b}$) to design the shoulder structure, and two of them ($d_{v}$ and $d_{b}$) an be adjusted without reconstructing the exoskeleton. A suitable set of parameters ($\phi$=15°, $d_{v}$=80mm, $d_{b}$=10mm) was selected through kinematic modeling and ROM simulations. Constrained but sufficient ROM was guaranteed in both sagittal and horizontal F/E for performing overhead work, as shown in Table 5, while the structure with a relatively small volume (22cm × 48cm × 15cm) is designed to reduce the possibility of collision. Since volume information for other devices was not found, no comparison was made. However, our exoskeleton volume is visually more compact than its counterparts. The discrepancies between the ROM in simulations and experiments may be due to the omission of friction and contact forces in the simulations.
We had a trade-off for the shoulder movements, limiting the abduction/adduction movements of the shoulder, which are not the focus of overhead work. Through the coordination of sagittal and horizontal F/E joints, the exoskeleton can achieve abduction/adduction movements in the anterior space of the body, but there are limitations on movements in the posterior area of the body, such as extending the elbow diagonally backward and outward. This method meets the design requirements of the overhead work, and similar degree of freedom configuration strategies can be found on the ShoulderX [47] and Airframe [22]. As shown in Fig. 13, wearing the exoskeleton reduced hand workspace, likely due to a corresponding reduction in shoulder ROM.
5.2 Comparison of Adaptability with Existing Exoskeletons
Various overhead tasks and individuals usually have different PATAs and thus different optimal torque profiles [25], [26]. Consequently, the task and individual adaptability of the exoskeleton are crucial, which essentially involves adjustable PATA and peak torque. Existing exoskeletons have made considerable efforts in adjusting the assistive levels with different peak torque. Several exoskeletons employ a manual knob or an electrical component to adjust the assistive level, such as the Shoulder X [21] and H-PULSE [9, 13], other devices require to replace the spring or other parts affecting the spring length. However, existing exoskeletons can only adjust the amplitude of peak torque, thus just adapting to a specific task. Employing torque profiles with a fixed PATA to cope with all different tasks and individuals is not an energy-efficient way and limits the adaptability of the torque generator [26], [31]. For such exoskeletons, if we need to adjust the PATA, we can only change it by reconstructing the exoskeleton.
To enhance the task adaptability and personalized customization capability of the exoskeleton, we proposed a torque generator with an adjustable PATA. We utilized a set of parameters ($\alpha$, $\beta$, $r_{1}$, $r_{2}$, $r_{3}$, $L_{i}$) to guide the design of the torque generator, and investigated their impacts on torque profiles via geometric analysis and simulations. A subset of parameters ($\alpha$, $\beta$, $L_{i}$) can be adjusted to reconfigure the exoskeleton, each controlling peak torque, PATA, and safety limit angle, respectively. Notably, $\beta$ can be adjusted within a range of -30º to 30º, resulting in a range of PATA from 90º to 150º, which meets the requirements of various overhead tasks and individuals. The experimental results demonstrated that monitored muscular activations in HIT-POSE Match condition decreased the most compared to the NO HIT-POSE condition, thus validating the effectiveness of PATA. However, the exoskeleton provided an assistive torque of around 2 N·m when lowering the arm, which may cause discomfort to users. This is inevitable in our current design and also a limitation of the current architecture. Although this is not reflected in users’ subjective evaluations, it may be attributed to the small sample size and individual sensitivity.
5.3 Effects of the Exoskeleton on Muscular Activations
For both tasks in HIT-POSE Match conditions, significantly reduced muscular activations were observed in agonist muscles (AD, MD, BB, PM, TR) and antagonist muscles (PD, LD) compared to NO HIT-POSE conditions. The exoskeleton provided support to the human shoulder, allowing agonist muscles to generate less force to achieve the same action, while also decreasing co-contraction of the antagonist muscles [13]. Similar results can be found in the ShoulderX [21], PAEXO [48], MATE [32], and H-VEX [26]. This phenomenon can be explained by the fact that there is a match between the PATA and the target task angle, resulting in increased assistive torque, with reductions consistent with the level of assistance [13, 49]. This also demonstrates the necessity of PATA for effective assistance. As reported by previous studies [50, 51], the reduced muscular activation implies lower force, resulting in decreased compressive force on the shoulder, and thus potentially reducing biomechanical stress on the shoulder girdle. Such positive effects suggest that long-term use of the exoskeleton has the potential to reduce the incidence of WMSDs, as reported by previous studies [17], [24].
Additionally, for screwing tasks in HIT-POSE Match conditions, the primary muscles (AD, MD, PD, BB, PM, and LD) significantly reduced their activations compared to HIT-POSE Mismatch conditions. The phenomenon is likely since a larger assistance was provided in HIT-POSE Match conditions, which provides evidence of our rational design. Moreover, AD muscle in screwing at a low height did not exhibit a significant decrease between the two conditions. A previous study has shown that the activation reduction of AD muscle in high tasks is greater than that in low tasks [11], which may be a possible reason. Furthermore, the muscular activations of BB and TR muscles in screwing at a high height did not significantly change between the two conditions. In screwing at a high height, we observed that most participants involved more arm extension (except for lowering the arm), compared to screwing at a low height. The mechanical hysteresis of the spring contributes to larger assistance involving arm extension, which may preferentially transmit to BB muscle.
Notably, compared to NO HIT-POSE conditions, reductions in ES muscle were also expected for both tasks in HIT-POSE Match conditions. This behavior is consistent with the existing study [52]. This may be because the exoskeleton transmits forces to the waist through the device’s back bar and waist belt. This result could indicate that HIT-POSE doesn’t increase the risk of back WMSDs. An unexpected phenomenon is that the muscular activations in Fig. 14 were quite high. This could be attributed to our different normalization method, which does not affect the trend of EMG changes.
5.4 Effects of The Exoskeleton on Subjective Feedback
For both tasks in HIT-POSE Match conditions, the RPE displayed statistical significance from the other two conditions. Compared to NO HIT-POSE conditions, scores of the physical demand and effort items exhibited a significant reduction for both tasks in HIT-POSE Match conditions. Participants perceived the greatest assistance and thus exerted the lowest physical effort. This outcome is in line with the existing study [13]. In addition, for both tasks in HIT-POSE Match conditions, scores of the frustration item also exhibited a significant reduction with NO HIT-POSE conditions, which is consistent with the previous study [24]. Evidence may be available in a SUS question (9. I felt very confident using the system.) with a high average score (4.56 ± 0.53). We speculated that mental confidence bridges the frustration from the task, which may be related to a placebo effect. Meanwhile, the SUS score (79.7 ± 5.9) expressed good usability of the exoskeleton according to the empirical evaluation [53].
5.5 Limitations of The Study
While the experimental evaluations have shown positive effects, this study still has some limitations. Firstly, the PATA of the torque generator needs to be regulated manually. Consequently, one of our future works is to add a lightweight electrical element to adjust PATA automatically. Secondly, this experiment was conducted in a laboratory setting. Accordingly, the participants had no work experience. A field study and even the female participants would be considered in our further research. Thirdly, the long-term effect of the exoskeleton on users is not studied, which may exhibit different results. As the proposed prototype matures, we will investigate this issue, aiming to get a more complete and accurate assessment.
6 CONCLUSION
In this study, we propose a passive occupational shoulder exoskeleton named HIT-POSE for overhead tasks, which enhances the adaptability of tasks and individuals while ensuring sufficient shoulder ROM. Firstly, we design an ergonomic shoulder structure by kinematic modeling and ROM simulations and verify through experiments. Secondly, we implement a torque generator with adjustable PATA through geometric analysis and simulations of torque profiles. Thirdly, we investigate the parameters’ impacts on torque profiles to determine adjustable parameters to achieve various torque profiles with different PATA without reconstructing the exoskeleton. Finally, we conduct experimental evaluations with aspects of objective and subjective within three conditions for two tasks, to validate the effectiveness of our proposed exoskeleton. Overall, the results have shown that HIT-POSE can decrease the physical load on users and has the potential to reduce the incidence of shoulder WMSDs.
In the future, we will strive to enhance the kinematic compliance of the exoskeleton to achieve a full ROM for the user. More realistic simulation algorithms of ROM will also be investigated to reduce the deviation between simulation and reality. In addition, a unique cam structure will be designed to further reduce the assistive torque at low elevation angles and improve the user acceptance and usability of the exoskeleton.
Acknowledgment
We gratefully acknowledge the participation of all individuals involved in this experiment. Their dedication and support were crucial to the completion of our research.